OpenStudy (-welp-):

"Simplify (-5-sqrt3)^2" How do I solve this?

OpenStudy (-welp-):

What did you do?

OpenStudy (mariamikayla):

(-5-rad(3))^2= (-5)^2 + 2* (-5)* (-rad3)+ (-rad(3))^2= 25+10rad3 +9= 34+10rad3

TheSmartOne (thesmartone):

@karim728 (a - b)^2 = a^2 - 2ab + b^2

OpenStudy (karim728):

oops yef i taught only rad 3 was being raised..

OpenStudy (karim728):

u gotta foil then as @TheSmartOne stated

OpenStudy (-welp-):

@TheSmartOne Thanks

TheSmartOne (thesmartone):

Anytime! :)